OpenStudy (anonymous):

Find volume or state it does not exist Region bounded by f(x) = sqrt((x+1)/(x^3)) and the x axis on the interval [1,infinite) is revolved about the x axis I though of using the shell method so I got this far 2pi integral from 1 to infinite x(sqrt((x+1)/(x^3))) Maybe the disk method would be easier...any thoughts? Either way I am stuck here.

OpenStudy (theviper):

@max0236 how did you writ this >>> \(\Huge{}\)

OpenStudy (anonymous):

\[2\pi \int\limits_{1}^{\infty} x \sqrt{(x+1)/x^3}dx\]

OpenStudy (anonymous):

try discs:

OpenStudy (anonymous):

you get a finite volume.

OpenStudy (anonymous):

questions?

OpenStudy (anonymous):

I finally got \[\frac{ 2b+1 }{ 2b^2 }- \frac{ 2(1)+1 }{ 2(1)^2 } \] now I substituted in b for infinite. So i plugged in b and 1 to the function. So how do I know its finite

OpenStudy (anonymous):

the limit is n-> infinity (-2 +1/n) / n aka zero :)

OpenStudy (anonymous):

whoops left out a '2' : bit off... the integral is - (2x+1)/2*x^2 so : { lim n->inf of -(2n+1)/n^2 } - - (2(1) +1)/(2*1^2) (all multiplied by pi)

OpenStudy (anonymous):

right a number over infinite goes to zero I recall now. Thank you very much!

OpenStudy (anonymous):

sure:)